# MultiGPrompt for Multi-Task Pre-Training and Prompting on Graphs

### 14 Dec 2023 v2

@@ 4.2. Multi-task Pre-training
In this part, we discuss the first stage on multi-task pre-training. In general, any graph-based pretext tasks can be used in our framework. Without loss of generality, in our experiments, we leverage there well-known pretext tasks, namely, DGI (Veličković et al., 2018), GraphCL (You et al., 2020), and link prediction (Liu et al., 2023a). We aim to aggregate the losses of multiple pretext tasks in a synergistic manner under the guidance of pretext tokens.
Pretext tokens. Assume we employ a total of $K$ pretext tasks for multi-task pre-training. Different pretext tasks focus on diverse aspects, and each has its unique loss function. Directly optimizing the sum of the $K$ losses leads to interference between tasks (Wu et al., 2020b; Yu et al., 2020; Wang et al., 2022b), and degrades the pre-training efficacy.
To avoid task interference, we leverage the concept of pretext tokens, which have been used to reformulate the task input in an earlier approach for pre-training language models (Wang et al., 2022b). In the context of graph, different layers of the graph encoder may have different emphasis on the representation, and thus carry variable significance to different pretext tasks. For instance, the input layer focuses on individual nodes’ features and thus are more important to node-level pretext tasks, while the hidden or output layers focus more on subgraph or graph features and thus are more important to local or graph-level tasks. Inspired by GraphPrompt+ (Yu et al., 2023b), we introduce a series of pretext tokens for each pretext task, to modify the input, hidden and output layers of the graph encoder alike.
@@
where $\{{\alpha}_{l}:0\leq l\leq L\}$ are hyperparameters. Typically, a graph encoder adopts a shallow architecture with few layers. In our implementation, we have $L=1$ and thus only two ${\alpha}_{l}$’s. Furthermore, as ${\alpha}_{l}$’s adjust the relative weights across layers, one of them can be fixed to 1, which means there is effectively only one free hyperparameter here.
@@
where $\{\beta_{k}:1\leq k\leq K\}$ contains $K$ hyperparameters, =$\mathcal{H}=\{\mathbf{H}_{\langle 1\rangle},\ldots,\mathbf{H}_{\langle K\rangle}\}$ represents the collection of task-specific embeddings, and $\mathcal{T}=\{\mathcal{T}_{\langle 1\rangle},\ldots,\mathcal{T}_{\langle K\rangle}\}$ denotes the collection of pretext token sets. The overall loss is optimized by updating the pretext tokens $\mathcal{T}$ and encoder weights $\Theta$. Note that the number of pretext tasks $K$ should be a small constant, with $K=3$ in our experiments. Similar to $\alpha_{l}$’s, one of $\beta_{k}$’s can be fixed at 1, leaving only two free hyperparameters.
Specifically, we employ DGI (Veličković et al., 2018), GraphCL (You et al., 2020), and Link Prediction (Liu et al., 2023a) as pretext tasks. Additional details are provided in Appendix D.
@@ 4.3. Prompting for Downstream Tasks
Open prompt. Similar to a composed prompt, an open prompt $\mathbf{p}_{{\langle\text{op}\rangle},l}\in\mathcal{P}_{\langle\text{op}\rangle}$ modifies the $l$-th layer of the pre-trained graph encoder. However, unlike the composed prompts, $\mathbf{p}_{{\langle\text{op}\rangle},l}$ is directly tuned, instead of being composed from the pretext tokens. In this way, an open prompt does not extract pre-trained knowledge specific to any pretext task, but focus on the global pre-trained model holistically.
@@
where $\Delta$ denotes the set of learnable parameters of the aggregation function.
@@ 5. Experiments
GCN 28.57 $\pm$ 5.07 31.27 $\pm$ 4.53 43.31 $\pm$ 9.35 48.08 $\pm$ 4.71 56.33 $\pm$ 10.40 50.95 $\pm$ 23.48 50.56 $\pm$ 3.01 17.10 $\pm$ 3.53
GAT 28.40 $\pm$ 6.25 30.76 $\pm$ 5.40 31.79 $\pm$ 20.11 35.32 $\pm$ 18.72 50.69 $\pm$ 23.66 50.58 $\pm$ 26.16 50.59 $\pm$ 12.43 16.80 $\pm$ 2.97
DGI/InfoGraph 54.11 $\pm$ 9.60 45.00 $\pm$ 9.19 45.22 $\pm$ 11.09 48.05 $\pm$ 14.83 52.57 $\pm$ 18.14 54.62 $\pm$ 15.36 48.21 $\pm$ 12.35 21.69 $\pm$ 5.98
GraphCL 51.96 $\pm$ 9.43 43.12 $\pm$ 9.61 46.15 $\pm$ 10.94 48.88 $\pm$ 15.98 54.11 $\pm$ 16.63 54.29 $\pm$ 17.31 53.69 $\pm$ 11.92 21.57 $\pm$ 5.20
GPPT 15.37 $\pm$ 4.51 21.45 $\pm$ 3.45 35.15 $\pm$ 11.40 35.37 $\pm$ 9.37 - - - -
GraphPrompt 54.25 $\pm$ 9.38 45.34 $\pm$ 10.53 47.22 $\pm$ 11.05 53.54 $\pm$ 15.46 54.60 $\pm$ 10.53 54.35 $\pm$ 14.78 54.73 $\pm$ 8.87 25.06 $\pm$ 7.56
MultiGPrompt 57.72 $\pm$ 9.94 54.74 $\pm$ 11.57 48.09 $\pm$ 11.49 54.47 $\pm$ 15.36 60.07 $\pm$ 12.48 56.17 $\pm$ 12.84 56.02 $\pm$ 8.27 26.63 $\pm$ 6.22
@@
In this section, we undertake comprehensive experiments across six benchmark datasets, to evaluate the efficacy of the proposed MultiGPrompt on few-shot node classification and graph classification tasks.
@@ 5.1. Experimental Setup
(3) Graph prompt-based learning: GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) are included under this umbrella. Their modus operandi revolves around leveraging link prediction during pre-training, and unifying downstream tasks into a common template as the pretext task. Both of them utilizes a single pretext task, and subsequently a single type of prompt in downstream adaptation. Note that GPPT is purposely designed to work with the downstream task of node classification, and cannot be directly applied to graph classification. Hence, in our experiments, we only employ GPPT for node classification.
@@ 5.2. Few-shot Performance Evaluation
One-shot node classification. The results are presented in Table 2. We make the following observations.
First, MultiGPrompt surpasses all baselines on all four datasets, indicating its advantage in the overall strategy of multi-task pre-training. We will further conduct a series of ablation studies in Sect. 5.3 to evaluate the importance of specific designs. Second, pre-training methods (DGI/InfoGraph, GraphCL) generally outperform supervised methods (GCN, GAT), as the former group leverage a pre-trained model. The results highlight the importance of acquiring general knowledge from label-free graphs. Lastly, “pre-train, prompt” methods, such as GraphPrompt and our MultiGPrompt, can further outperform the pre-training approaches without prompts, demonstrating the advantage of prompt-based learning epsecially in few-shot settings.
@@ 5.3. Ablation study
We start with three basic variants that only utilize a single pretext task: using only DGI/InfoGraph (DGI), GraphCL, and link prediction (LP). These three basic variants simply employ a classifier during downstream fine-tuning, without any prompting. We further compare three more advanced variants, namely, DGI+, GraphCL+ and LP+, which has the exact same architecture and dual-prompt design as the full model MultiGPrompt, but only utilize one pretext task. Referring to Fig. 5, we observe that MultiGPrompt consistently outperforms all variants using a single pretext task, with or without prompts. This finding underscores the value of leveraging multiple pretext tasks.
@@ A. Further Descriptions of Datasets
(1) Cora [^4] https://relational.fit.cvut.cz/dataset/CORA (McCallum et al., 2000) comprises 2,708 scientific publications, each categorized into one of seven classes. The citation network encompasses 5,429 links. Every publication in the dataset is depicted by a 0/1-valued word vector, indicating the absence/presence of the corresponding word from the dictionary, which consists of 1433 unique words.
(2) Citeseer [^5] https://nrvis.com/download/data/labeled/citeseer.zip (Sen et al., 2008) is composed of 3,312 scientific publications, each categorized into one of six classes. The citation network entails 4,732 links. Each publication in the dataset is characterized by a 0/1-valued word vector, indicating the absence/presence of the corresponding word from the dictionary, which encompasses 3,703 unique words.
(3) PROTEINS [^6] https://www.chrsmrrs.com/graphkerneldatasets/PROTEINS.zip (Borgwardt et al., 2005) encompasses a collection of protein graphs, embodying various attributes including the amino acid sequence, conformation, structure, and distinctive features such as active sites of the proteins. In this collection, each node signifies the secondary structures, whereas each edge represents the neighboring relation either within the amino-acid sequence or in 3D space. The nodes are classified into three categories, while the graphs are divided into two classes.
(4) ENZYMES [^7] http://www.chrsmrrs.com/graphkerneldatasets/ENZYMES.zip (Wang et al., 2022a) constitutes a dataset comprising 600 enzymes, meticulously collected from the BRENDA enzyme database. These enzymes are meticulously categorized into 6 distinct classes in accordance with their top-level EC enzyme classification.
(5) BZR [^8] https://www.chrsmrrs.com/graphkerneldatasets/BZR.zip (Rossi and Ahmed, 2015) encompasses a collection of 405 ligands, each associated with the benzodiazepine receptor and graphically represented as individual entities. The entire ligand set is bifurcated into 2 distinct categories.
(6) COX2999https://www.chrsmrrs.com/graphkerneldatasets/COX2.zip (Rossi and Ahmed, 2015) encompasses a dataset that includes 467 molecular structures, specifically of cyclooxygenase-2 inhibitors, wherein each node symbolizes an atom and each edge represents the chemical bond—be it single, double, triple, or aromatic—between atoms. The entirety of the molecules is categorized into two classes.
@@ B. Further Descriptions of Baselines
In this section, we present more details for the baselines.
@@
  • GCN (Kipf and Welling, 2017): GCN employs a mean-pooling-based neighborhood aggregation approach to consolidate messages from adjacent nodes.
  • GAT (Veličković et al., 2018): GAT also leverages neighborhood aggregation for end-to-end node representation learning, distinguishes itself by allocating varied weights to neighboring nodes, thus modifying their influence in the aggregation process.
@@
  • DGI (Veličković et al., 2018): DGI operates as a self-supervised pre-training methodology tailored for homogeneous graphs. It is predicated on the maximization of mutual information (MI), aiming to enhance the estimated MI between locally augmented instances and their global counterparts.
  • InfoGraph (Sun et al., 2019): Expanding upon DGI, InfoGraph is centered on graph-level tasks, endeavoring to maximize the alignment between node and graph embeddings.
  • GraphCL (You et al., 2020): GraphCL leverages a variety of graph augmentations for self-supervised learning, tapping into the intrinsic structural patterns of graphs. The overarching goal is to amplify the concordance between different augmentations throughout graph pre-training.
@@
  • GPPT (Sun et al., 2022): GPPT embraces a GNN model, pre-trained through executing a link prediction task. The utilization of a prompt module structures the downstream node classification task, orchestrating it cohesively with the link prediction format.
  • GraphPrompt (Liu et al., 2023a): GraphPrompt employs subgraph similarity calculations as a mechanism to amalgamate pre-training and downstream tasks, inclusive of node and graph classification. A learnable prompt is subsequently refined during the execution of the downstream task to incorporate task-specific nuances.
C. Implementation Details of Approaches
@@
For DGI (Veličković et al., 2018), we utilize a 1-layer GCN as the base model and set the hidden dimension to 256. Additionally, we employ prelu as the activation function. For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its base model, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2.
For GPPT (Sun et al., 2022), we utilize a 2-layer GraphSAGE as its base model, setting the hidden dimension to 256. For base GraphSAGE, we also employ a mean aggregator. For GraphPrompt (Liu et al., 2023a), a 3-layer GCN is used as the base model for all datasets, with the hidden dimensions set to 256.
Details of MultiGPrompt. For our proposed MultiGPrompt, we utilize a 1-layer GCN as the base model for all datasets, assigning the hidden dimensions a value of 256. We designate $\alpha_{0}=0.0001$ for node classification tasks, while setting $\alpha_{0}=0$ for graph classification tasks. And $\alpha_{1}$ is set to $1$. The parameters $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are set to $0.9$, $0.9$, and $0.1$ respectively.
@@ D. Pretext Tasks
In our experiments, we employ three novel pretext tasks, i.e., DGI (Veličković et al., 2018), GraphCL (You et al., 2020) and link prediction (Liu et al., 2023a). Consequently, given a pretext task $k$ and graph $G$. $\mathcal{T}_{\langle k\rangle}$ serves as the pretext tokens for pretext task $k$. $\mathcal{A}$ denotes the set of positive samples, $\mathcal{B}$ represents the set of corresponding negative samples. Define $\vec{h}_{{\langle k\rangle},v}$, a row of $\vec{H}_{\langle k\rangle}$, as node $v$’s representation. $\vec{h}_{\langle k\rangle,G}$ denotes the representation of $G$, that is:
@@
Therefore, DGI’s loss function $\mathcal{L}_{\text{pre}_{\langle\text{DGI}\rangle}}(\vec{H}_{{\langle\text{DGI}\rangle}};\mathcal{T}_{\langle\text{DGI}\rangle},\Theta)=$
(15) $\frac{1}{|\mathcal{A}|+|\mathcal{B}|}(\sum_{a\in\mathcal{A}}\log[\text{sim}(\vec{h}_{\langle\text{DGI}\rangle,a},\vec{h}_{{\langle\text{DGI}\rangle},G})]+\\ \sum_{b\in\mathcal{B}}\log[1-\text{sim}(\vec{h}_{{\langle\text{DGI}\rangle},b},\vec{h}_{{\langle\text{DGI}\rangle},G})]),\end{aligned}$
where $|\mathcal{A}|$ represents the number of positive samples, similarly $|\mathcal{B}|$ serves as the number of negative samples.
Similarly, the pre-training loss for GraphCL (hereinafter abbreviated as GCL) $\mathcal{L}_{\text{pre}_{\langle\text{GCL}\rangle}}(\vec{H}_{{\langle\text{GCL}\rangle}};\mathcal{T}_{\langle\text{GCL}\rangle},\Theta)=$
(16) $-\ln\frac{\exp(\text{sim}(\mathbf{h}_{\langle\text{GCL}\rangle,a},\mathbf{h}_{{\langle\text{GCL}\rangle},v})/\tau)}{\sum_{b\in\mathcal{B}}\exp(\text{sim}(\mathbf{h}_{{\langle\text{GCL}\rangle},b},\mathbf{h}_{\langle\text{GCL}\rangle,v})/\tau)}.$
Finally, let the $1$-hop subgraph of node $v$ be denoted as $S_{v}$, and the pre-training loss for link prediction (subsequently abbreviated as LP) is $\mathcal{L}_{\text{pre}_{\langle\text{LP}\rangle}}(\vec{H}_{{\langle\text{DGI}\rangle}};\mathcal{T}_{\langle\text{LP}\rangle},\Theta)=$
@@
where $\vec{h}_{{\langle\text{LP}\rangle},S_{v}}$ serves as the vector representation of $S_{v}$ with pretext tokens $\mathcal{T}_{\langle\text{LP}\rangle}$ which is calculated in the same way as Eq. (14). $\mathcal{X}_{\text{pre}}$ denotes the sampled tuples utilized for link prediction (Liu et al., 2023a), $a$ serves as positive sample and $b$ denotes negative samples.

### 22 Feb 2024 v6

@@ 4.2. Multi-task Pre-training
In this part, we discuss the first stage on multi-task pre-training. In general, any graph-based pretext tasks can be used in our framework. In our experiments, we leverage three well-known pretext tasks, namely, DGI (Veličković et al., 2018), GraphCL (You et al., 2020), and link prediction (Yu et al., 2023b). We aim to aggregate the losses of pretext tasks in a synergistic manner under the guidance of pretext tokens.
Pretext tokens. Assume we employ a total of $K$ pretext tasks for multi-task pre-training. Different pretext tasks focus on diverse aspects, and each has its unique loss function. Directly optimizing the sum of the $K$ losses leads to interference between tasks (Wu et al., 2020b; Yu et al., 2020; Wang et al., 2022c), and degrades the pre-training efficacy.
To avoid task interference, we leverage the concept of pretext tokens, which have been used to reformulate the task input in an earlier approach for pre-training language models (Wang et al., 2022c). In the context of graph, different layers of the graph encoder may have different emphasis on the representation, and thus carry variable significance to different pretext tasks. For instance, the input layer focuses on individual nodes’ features and thus are more important to node-level pretext tasks, while the hidden or output layers focus more on subgraph or graph features and thus are more important to local or graph-level tasks. Inspired by GraphPrompt+ (Yu et al., 2023b), for each pretext task, we apply a series of pretext tokens to modify the input, hidden and output layers of the graph encoder alike.
@@
where $\{{\alpha}_{l}:0\leq l\leq L\}$ are hyperparameters.
@@
where $\{\beta_{k}:1\leq k\leq K\}$ contains $K$ hyperparameters, =$\mathcal{H}=\{\mathbf{H}_{\langle 1\rangle},\ldots,\mathbf{H}_{\langle K\rangle}\}$ represents the collection of task-specific embeddings, and $\mathcal{T}=\{\mathcal{T}_{\langle 1\rangle},\ldots,\mathcal{T}_{\langle K\rangle}\}$ denotes the collection of pretext token sets. The overall loss is optimized by updating the pretext tokens $\mathcal{T}$ and encoder weights $\Theta$. Note that the number of pretext tasks $K$ should be a small constant, with $K=3$ in our experiments.
@@ 4.3. Prompting for Downstream Tasks
Open prompt. Similar to a composed prompt, an open prompt $\mathbf{p}_{{\langle\text{op}\rangle},l}\in\mathcal{P}_{\langle\text{op}\rangle}$ modifies the $l$-th layer of the pre-trained graph encoder. Unlike the composed prompts, $\mathbf{p}_{{\langle\text{op}\rangle},l}$ is directly tuned, instead of being composed from the pretext tokens. In this way, an open prompt does not extract pre-trained knowledge specific to any pretext task, but focus on the global pre-trained model holistically.
@@
where $\Delta$ denotes learnable parameters of the aggregation function.
@@ 5. Experiments
GCN 28.57 $\pm$ 5.07 31.27 $\pm$ 04.53 43.31 $\pm$ 09.35 48.08 $\pm$ 04.71 56.33 $\pm$ 10.40 50.95 $\pm$ 23.48 50.56 $\pm$ 03.01 17.10 $\pm$ 3.53
GAT 28.40 $\pm$ 6.25 30.76 $\pm$ 05.40 31.79 $\pm$ 20.11 35.32 $\pm$ 18.72 50.69 $\pm$ 23.66 50.58 $\pm$ 26.16 50.59 $\pm$ 12.43 16.80 $\pm$ 2.97
DGI/InfoGraph 54.11 $\pm$ 9.60 45.00 $\pm$ 09.19 45.22 $\pm$ 11.09 48.05 $\pm$ 14.83 52.57 $\pm$ 18.14 54.62 $\pm$ 15.36 48.21 $\pm$ 12.35 21.69 $\pm$ 5.98
GraphCL 51.96 $\pm$ 9.43 43.12 $\pm$ 09.61 46.15 $\pm$ 10.94 48.88 $\pm$ 15.98 54.11 $\pm$ 16.63 54.29 $\pm$ 17.31 53.69 $\pm$ 11.92 21.57 $\pm$ 5.20
GPPT 15.37 $\pm$ 4.51 21.45 $\pm$ 03.45 35.15 $\pm$ 11.40 35.37 $\pm$ 09.37 - - - -
GraphPrompt 54.25 $\pm$ 9.38 45.34 $\pm$ 10.53 47.22 $\pm$ 11.05 53.54 $\pm$ 15.46 54.60 $\pm$ 10.53 54.35 $\pm$ 14.78 54.73 $\pm$ 08.87 25.06 $\pm$ 7.56
MultiGPrompt 57.72 $\pm$ 9.94 54.74 $\pm$ 11.57 48.09 $\pm$ 11.49 54.47 $\pm$ 15.36 60.07 $\pm$ 12.48 56.17 $\pm$ 12.84 56.02 $\pm$ 08.27 26.63 $\pm$ 6.22
@@
In this section, we undertake comprehensive experiments across six benchmark datasets, to evaluate the efficacy of the proposed MultiGPrompt on few-shot node and graph classification tasks. The datasets and code are available. [^2] https://github.com/Nashchou/MultiGPrompt
@@ 5.1. Experimental Setup
(3) Graph prompt-based learning: GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) are included under this umbrella. They first leverage link prediction during pre-training, and unify downstream tasks into a common template as the pretext task. Both of them utilizes a single pretext task, and subsequently a single type of prompt in downstream adaptation. Note that GPPT is purposely designed to work with the downstream task of node classification, and cannot be directly applied to graph classification. Hence, in our experiments, we only employ GPPT for node classification.
@@ 5.2. Few-shot Performance Evaluation
One-shot node classification. The results are presented in Table 2. We make the following observations. First, MultiGPrompt surpasses all baselines on all four datasets, indicating its advantage in the overall strategy of multi-task pre-training. We will further conduct a series of ablation studies in Sect. 5.3 to evaluate the importance of specific designs. Second, pre-training methods (DGI/InfoGraph, GraphCL) generally outperform supervised methods (GCN, GAT), as the former group leverage a pre-trained model, which highlights the importance of acquiring general knowledge from label-free graphs. Lastly, “pre-train, prompt” methods, such as GraphPrompt and our MultiGPrompt, can further outperform the pre-training approaches without prompts, demonstrating the advantage of prompt-based learning epsecially in few-shot settings.
@@ 5.3. Ablation study
We start with three basic variants that only utilize a single pretext task: using only DGI/InfoGraph (DGI), GraphCL, and link prediction (LP). These three basic variants employ a classifier during downstream fine-tuning, without prompting. We further compare three more advanced variants, namely, DGI+, GraphCL+ and LP+, which has the exact same architecture and dual-prompt design as the full model MultiGPrompt, but only utilize one pretext task. Referring to Fig. 5, we observe that MultiGPrompt consistently outperforms all variants using a single pretext task, with or without prompts, which underscores the value of leveraging multiple pretext tasks.
@@ 6. Conclusions
Acknowledgments
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore. This work is also supported in part by the National Key Research and Development Program of China under Grant 2020YFB2103803.
@@ A. Further Descriptions of Datasets
(1) Cora [^5] https://paperswithcode.com/dataset/cora (McCallum et al., 2000) comprises 2,708 scientific publications, each categorized into one of seven classes. The citation network encompasses 5,429 links. Every publication in the dataset is depicted by a 0/1-valued word vector, indicating the absence/presence of the corresponding word from the dictionary, which consists of 1433 unique words.
(2) Citeseer [^6] https://paperswithcode.com/sota/node-classification-on-citeseer (Sen et al., 2008) is composed of 3,312 scientific publications, each categorized into one of six classes. The citation network entails 4,732 links. Each publication in the dataset is characterized by a 0/1-valued word vector, indicating the absence/presence of the corresponding word from the dictionary, which encompasses 3,703 unique words.
(3) PROTEINS [^7] https://chrsmrrs.github.io/datasets/docs/datasets/ (Borgwardt et al., 2005) encompasses a collection of protein graphs, embodying various attributes including the amino acid sequence, conformation, structure, and distinctive features such as active sites of the proteins. In this collection, each node signifies the secondary structures, whereas each edge represents the neighboring relation either within the amino-acid sequence or in 3D space. The nodes are classified into three categories, while the graphs are divided into two classes.
[…]
@@ B. Further Descriptions of Baselines
In this section, we present more details for the baselines used in our experiments.
@@
  • GCN (Kipf and Welling, 2017): GCN employs a mean-pooling-based neighborhood aggregation approach to process messages from adjacent nodes.
  • GAT (Veličković et al., 2018): GAT also leverages neighborhood aggregation for end-to-end node representation learning. However, during the aggregation process, it allocates varied weights to neighboring nodes through an attention mechanism.
@@
  • DGI (Veličković et al., 2018): DGI operates as a self-supervised pre-training method. It is predicated on the maximization of mutual information, aiming to enhance the consistency between locally augmented instances and their global counterparts.
  • InfoGraph (Sun et al., 2019): Expanding upon DGI, InfoGraph is centered on graph-level tasks, aiming to maximize the alignment between node and graph embeddings.
  • GraphCL (You et al., 2020): GraphCL leverages a variety of graph augmentations for self-supervised learning, tapping into various intrinsic properties of the graphs for pre-training.
@@
  • GPPT (Sun et al., 2022): GPPT employs a GNN model, pre-trained on a link prediction task. A prompt module is designed to facilitate the downstream node classification task, which is converted to a link prediction format.
  • GraphPrompt (Liu et al., 2023a): GraphPrompt employs subgraph similarity as a mechanism to unify different pretext and downstream tasks, including link prediction, node classification, and graph classification. A learnable prompt is subsequently tuned for each downstream task.
C. Implementation Details
@@
For DGI (Veličković et al., 2018), we utilize a 1-layer GCN as the base model and set the hidden dimensions to 256. Additionally, we employ prelu as the activation function. For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its base model, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2.
For GPPT (Sun et al., 2022), we utilize a 2-layer GraphSAGE as its base model, setting the hidden dimensions to 256. For base GraphSAGE, we also employ a mean aggregator. For GraphPrompt (Liu et al., 2023a), a 3-layer GCN is used as the base model for all datasets, with the hidden dimensions set to 256.
Details of MultiGPrompt. For our proposed MultiGPrompt, we utilize a 1-layer GCN as the base model for all datasets, with the hidden dimensions set to 256. We set $\alpha_{0}=0.0001$ for node classification tasks, while setting $\alpha_{0}=0$ for graph classification tasks. Meanwhile, $\alpha_{1}$ is set to $1$ for both tasks. The parameters $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are set to $0.9$, $0.9$, and $0.1$, respectively.
@@ D. Pretext Tasks
In our experiments, we employ three pretext tasks, i.e., DGI (Veličković et al., 2018), GraphCL (You et al., 2020) and link prediction (Liu et al., 2023a). Consider a pretext task $k$ and a graph $G$. $\mathcal{T}_{\langle k\rangle}$ serves as the pretext tokens for the task $k$. Define $\vec{h}_{{\langle k\rangle},v}$, a row of $\vec{H}_{\langle k\rangle}$, as node $v$’s representation specific to task $k$. Further let $\vec{h}_{\langle k\rangle,G}$ denote the representation of $G$ for task $k$, which is given by
@@
Then, the loss function of DGI can be expressed as follows.
(15) $\sum_{G\in\mathcal{X}_{\text{pre}}}\frac{1}{|\mathcal{A}_{G}|+|\mathcal{B}_{G}|}(\sum_{a\in\mathcal{A}_{G}}\log[\text{sim}(\vec{h}_{\langle\text{DGI}\rangle,a},\vec{h}_{{\langle\text{DGI}\rangle},G})]+\\ \sum_{b\in\mathcal{B}_{G}}\log[1-\text{sim}(\vec{h}_{{\langle\text{DGI}\rangle},b},\vec{h}_{{\langle\text{DGI}\rangle},G})]),\end{aligned}$
where $\mathcal{X}_{\text{pre}}$ is a set of pre-training graphs, and $\mathcal{A}_{G},\mathcal{B}_{G}$ represent the set of positive and negative instances for $G$, respectively.
Similarly, the pre-training loss for GraphCL (abbreviated as GCL) can be formulated as follows.
(16) $-\sum_{G\in\mathcal{X}_{\text{pre}}}\ln\frac{\exp(\text{sim}(\mathbf{h}_{\langle\text{GCL}\rangle,a\in\mathcal{A}_{G}},\mathbf{h}_{{\langle\text{GCL}\rangle},v\in V_{G}})/\tau)}{\sum_{b\in\mathcal{B}_{G}}\exp(\text{sim}(\mathbf{h}_{{\langle\text{GCL}\rangle},b},\mathbf{h}_{\langle\text{GCL}\rangle,v\in V_{G}})/\tau)},$
where $V_{G}$ denotes the set of nodes in graph $G$, and $\tau$ is a temperature hyperparameter.
Finally, let the $1$-hop subgraph of node $v$ be denoted as $S_{v}$. Then, the pre-training loss for link prediction (abbreviated as LP) can be formulated as follows.
@@
where $\vec{h}_{{\langle\text{LP}\rangle},S_{v}}$ serves as the vector representation of $S_{v}$, which is calculated through a readout similar to Eq. (14). $\mathcal{X}_{\text{pre}}$ denotes the set of tuples sampled for link prediction (Liu et al., 2023a), where a tuple $(v,a,b)$ implies that $(v,a)$ is an edge (i.e., a positive example) and $(v,b)$ is not an edge (i.e., a negative example).